OpenStudy (anonymous):

Alfred graphed two lines, y = x + 4 and y = x - 2. How would this be graphed?

OpenStudy (anonymous):



OpenStudy (anonymous):

With a pen and paper.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

very funny xD

OpenStudy (anonymous):

thank you jose

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

just determine the slope of each line, (the rise/run)

OpenStudy (anonymous):

Based on the slope-intercept formula for a line: y=mx+b y=x+4 has a slope = 1 and a y-intercept = 4 y=x-2 has slope = 1 and y-intercept = -2 Write the slopes in the rise/run form. Each slope is 1/1 Graph the y-intercepts and from each of these points rise 1 unit and run to the right one unit. These locations are other points on the lines. Connect the points with a straightedge and graph the lines.